OpenStudy (anonymous):

Can someone please Help me with college algebra 2 ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sqrt[4]{32y^13}\]

OpenStudy (anonymous):

32 = 2 x 2 x 2 x 2

OpenStudy (anonymous):

i left out one 2

OpenStudy (anonymous):

its y raised to the 13th

OpenStudy (anonymous):

im guessing im factoring it?

OpenStudy (anonymous):

are you trying to solve for y?

OpenStudy (anonymous):

no i think you are trying to simplify the expression

OpenStudy (anonymous):

i was using prime factorization

OpenStudy (anonymous):

right simplify

OpenStudy (anonymous):

so you have 32 factored to 2^5, so take away 16 from inside the radical and leave in one 2 as for y, you have y^3 outside the radical and one y inside, since (y^3)^4 * y = y^13

OpenStudy (anonymous):

im pretty sure it is 2y^3*radical4(2y)

OpenStudy (anonymous):

so its 32^(1/4)*y^(13/4)*3^(1/4)

OpenStudy (anonymous):

its a 4 outside the radical and 32y^13 on the inside

OpenStudy (anonymous):

Factor completely: 6b^2+4a-8b-3ab

OpenStudy (anonymous):

wait, ok your previous expression was messed up, can you re type it?